OpenStudy (anonymous):

Find the intercepts in y =x^2 +3x/(3x+1)^2

OpenStudy (anonymous):

I found the y-intercept. I just need help with finding the xintercept because Im stuck and I dont know what to do next

OpenStudy (anonymous):

If you want to know how far I got, I'll type it out. I'm not sure if its correct though

OpenStudy (zale101):

\[0 =(x^2+3x)/(3x+1)^2 \] remove the fraction first and then multiply the sides by the denominator.

OpenStudy (anonymous):

So then you have x^2+3x= (3x+1)^2

OpenStudy (anonymous):

Then what do you do?

OpenStudy (zale101):

0 * (3x + 1)^2 = (x^2 + 3x)= 0*[(3x+1)^2] = 0 You can multiply out the denominator, as long as the denominator is not zero, get it?

OpenStudy (chenping):

basically juz cancel out the denominator right? so x=0 or x=-3 ? That's all? @@